OpenStudy (lerenstudy):

I'm not sure if I got the answer correct for this or not? Help please? The function f(x) = 582(7)^x represents the growth of a mosquito population every year in a remote swamp. Troy wants to manipulate the formula to an equivalent form that calculates every 2 months, not every year. Which function is correct for Troy's purposes?

OpenStudy (lerenstudy):

Answer choices are: \[f(x)=582(7^{1/6})^{6x}\] \[f(x)=52(7)^{x}\] \[f(x)=582(7^{12})^{x/12}\] \[f(x)=5852(7)^x\]

OpenStudy (lerenstudy):

I think it's the last one but I'm not sure

OpenStudy (anonymous):

Well, two months is \(1/6\)th of a year.

OpenStudy (lerenstudy):

right, so then the first one would be true?

OpenStudy (anonymous):

Yes

OpenStudy (lerenstudy):

And then the x/12 would work because you could replace x with 2 correct?

OpenStudy (anonymous):

Wait, let me think about it for a second

OpenStudy (anonymous):

Are you sure those are the answers?

OpenStudy (lerenstudy):

Yes, I just checked them over now

OpenStudy (anonymous):

The first one is the one that seems to make the most sense, I guess

OpenStudy (lerenstudy):

Right, but I have to eliminate all possible answers and find the one that doesn't work

OpenStudy (lerenstudy):

And the most random seems to be the final one

OpenStudy (seratul):

Maybe this could help. http://openstudy.com/updates/52ab9f5ae4b0b09acc7ee460

OpenStudy (lerenstudy):

^I looked at that but they're saying which one works, I think. I need the one that doesn't

OpenStudy (lerenstudy):

My logic right now on picking the last one is that the starting number wouldn't be more, would it?

OpenStudy (lerenstudy):

oh is CORRECT for the situation. That's such a weird phrasing. So then it would be the first one

OpenStudy (anonymous):

You should pick the ans. that will give the same results some length of time, say a year. so \[582*7^1 = ?\]

OpenStudy (anonymous):

can also check results for two months \[582*7^{(1/6)} = ??\]

OpenStudy (lerenstudy):

So the first would be 4109 and the second would be approx 805

OpenStudy (anonymous):

compare the one year formulation to the 2 month one \[582*(7^{1/6})^6 = 582*7^1\]

OpenStudy (anonymous):

six two months = 1 year

OpenStudy (anonymous):

wait! that doesn't make sense

OpenStudy (lerenstudy):

Right, they're not the same and that wouldn't work out, so would it be the third one, because otherwise you're just making it into a year again, right?

OpenStudy (anonymous):

the first and third are the same... for time in units of two months you would want \[582*(7^{1/6})^t\] which isn't an option?

OpenStudy (lerenstudy):

It's not

OpenStudy (lerenstudy):

I think maybe the x/12 wasn't an exponent. It's really hard to tell from what they gave me

OpenStudy (anonymous):

None of the others work... humm.

OpenStudy (anonymous):

you would need the x in an exponent.

OpenStudy (lerenstudy):

So then the first, because otherwise they would be the same?

OpenStudy (anonymous):

\[582*(7^{1/6})^{6x}=582*7^{6x/6}=582*7^x\]

OpenStudy (anonymous):

so the units x is in are still years.

OpenStudy (anonymous):

So idk what they want.

OpenStudy (lerenstudy):

right but just saying 2/12th of the final answer would be wrong, right?

OpenStudy (lerenstudy):

I'll just go through with the first and see if it comes out right or wrong

OpenStudy (anonymous):

it looks like the best bet.

OpenStudy (lerenstudy):

It said it was right, so I guess that works out

OpenStudy (anonymous):

I still don't buy it. but good it accepted it. :)

OpenStudy (lerenstudy):

Yup, thanks a lot still. I mean, that whole question was worded weird but we somehow managed to at least get it right

OpenStudy (anonymous):

np.